# Measurement as Bricolage: Examining How Data Scientists Construct Target Variables for Predictive Modeling Tasks

### v2

@@ 4.2.2. Composing
Yet some participants found that composing had simplicity limitations. For example, P2 tried to use a multi-outcome latent variable to predict student “engagement.” However, when P2 presented the model to software engineers and product managers, they distrusted its predictions when the model contradicted their prior beliefs about new product features. P2 explained that:
“If I show them a latent model, and it says, ‘No, it doesn’t look like you improved engagement’ then there’s gonna be a lot of questions about that model. What is this actually measuring? How do you explain it?”
@@ 5.1.1. The Data Science Repertoire
Materials in data scientists’ repertoire function as what Lévi-Strauss calls “signs”, or concrete objects that represent abstract concepts, but remain constrained by their imagined uses. P6’s creative re-purposing of raw user interactions in an online learning platform as a proxy for “reading ability” exemplifies the use of a sign, by unexpectedly transforming a mundane data field into an operationalization of a construct. Yet, as Levi-Strauss (1966) notes, “the possibilities remain always limited by the particular history of each piece, and by what is predetermined in it due to the original usage for which it was conceived.” Had P6 not been aware of this proxy for “reading ability”, established through a chance encounter with a colleague’s work, they may have never made this connection.
@@ 5.1.2. Engaging in Dialogue with Data
After taking stock of the repertoire, the bricoleur engages in dialogue with their materials to identify a workable solution to the problem at hand. As Lévi-Strauss explains, the bricoleur must “engage into a kind of dialog with [the repertoire], to index, before choosing among them, the possible answers that the set can offer to his problem.” (Levi-Strauss, 1966) This dialogue not only shapes the bricoleur’s understanding of each element in isolation, but also their interactions: “one element’s possibilities interact with all other elements’ possibilities, with the overall organization of the artifact he makes. The results of these interactions are never what he expects, and he must respond to them.” (Louridas, 1999) P10’s experience illustrates how dialogue unfolds in data science. After initially selecting a cardiovascular outcome based on precedent in academic literature (piggybacking), “seeing the data, getting to work with data” led them to swap to a renal outcome that was more predictable given available biomarkers. This example illustrates problem (re)formulation strategies serve as the brushstroke of measurement in data science work. Like a painter with a canvas, data scientists apply a strategy, step back to survey their progress along multiple criteria, then consider a new strategy in response to newly-identified flaws.
@@ 5.2. Design Implications: Reconciling Top-Down Measurement Interventions with Bricolage Practice
Our findings uncover a tension between existing interventions designed to help analysts implement the traditional top-down measurement workflow, and data scientists’ bottom-up bricolage practices. However, this tension is not unique to data science. History reveals similar patterns across domains where top-down planning approaches have failed while bottom-up alternatives succeeded. In urban planning, modernist urban renewal projects that imposed regimented public works projects onto complex social systems often destroyed functioning neighborhoods, while community-based planning approaches showed strong success (Fuller and Moore, 2017; Scott, 2020). In software development, rigid and systematic waterfall planning methodologies gave way to agile approaches that embrace iteration and adaptation (Osorio et al., 2011). Key to the success of these bottom-up alternatives was not wholesale rejection of structure, but rather targeted interventions like mixed-use zoning in urban planning and sprint retrospectives in agile development. These interventions recognized that while existing practices were in need of reform, effective solutions must work within resource constraints and pinpoint specific procedural weaknesses in need of constructive redirection. Similarly, when improving measurement practices in data science, interventions tailored to support data scientists’ existing practices are likely to be more effective than direct translation of rigid, top-down measurement frameworks. In the following sections, we examine how existing top-down measurement interventions might be adapted to better support data scientists’ bricolage practices.
@@ 5.3.3. Helping Data Scientists Apply Strategies
Refining. While refining, data scientists make granular adjustments to how a target variable is coded in their data. Participants often engaged in refining by modifying a label definition used to operationalize a construct using a set of indicators used to categorize instances. For instance, P13 engaged in refining by working with domain experts to spot gaps in label definitions, then patch them with a revised set of indicators. Though refining can improve validity, practitioners lack tools to encode domain expert’s knowledge into label definitions. Instead, existing tools developed by the HCI and ML communities help annotators apply a fixed label definition to annotate data (e.g., (Nakayama et al., 2018; Muller et al., 2021; Desmond et al., 2021; Daudert, 2020)) or use crowdsourcing to refine annotation guidelines (e.g., (Chen and Zhang, 2023; Bragg et al., 2018; Chang et al., 2017; K. Chaithanya Manam et al., 2019; Manam and Quinn, 2018)). [^4] Traditional annotation guidelines are de facto label definitions if they serve as a measurement instrument that operationalizes a latent construct of interest. Future HCI and ML research can help data scientists encode domain experts’ knowledge in label definitions by developing tools that help teams collaboratively systematize concepts (Adcock and Collier, 2001) by iteratively identifying and refining indicators used in label definitions. Because our results indicate that domain experts require rich contextual information to identify areas where the systemization of a concept is incomplete, tools supporting this process should capture rich contextual information (e.g., written notes) in data.

### v3

@@ 4.2.2. Composing
Yet some participants found that composing had simplicity limitations. For example, P2 tried to use a multi-outcome latent variable model to predict student “engagement.” However, when P2 presented the model to software engineers and product managers, they distrusted its predictions when the model contradicted their prior beliefs about new product features. P2 explained that:
“If I show them a latent [variable] model, and it says, ‘No, it doesn’t look like you improved engagement’ then there’s gonna be a lot of questions about that model. What is this actually measuring? How do you explain it?”
@@ 5.1.1. The Data Science Repertoire
Materials in data scientists’ repertoire function as what Lévi-Strauss calls “signs”, or concrete objects that represent abstract concepts, but remain constrained by their imagined uses. P6’s creative re-purposing of raw user interactions in an online learning platform as a proxy for “reading ability” exemplifies the use of a sign, by unexpectedly transforming a mundane data field into an operationalization of a construct. Yet, as Lévi-Strauss (1966) notes, “the possibilities remain always limited by the particular history of each piece, and by what is predetermined in it due to the original usage for which it was conceived.” Had P6 not been aware of this proxy for “reading ability”, established through a chance encounter with a colleague’s work, they may have never made this connection.
@@ 5.1.2. Engaging in Dialogue with Data
After taking stock of the repertoire, the bricoleur engages in dialogue with their materials to identify a workable solution to the problem at hand. As Lévi-Strauss explains, the bricoleur must “engage into a kind of dialog with [the repertoire], to index, before choosing among them, the possible answers that the set can offer to his problem.” (Lévi-Strauss, 1966) This dialogue not only shapes the bricoleur’s understanding of each element in isolation, but also their interactions: “one element’s possibilities interact with all other elements’ possibilities, with the overall organization of the artifact he makes. The results of these interactions are never what he expects, and he must respond to them.” (Louridas, 1999) P10’s experience illustrates how dialogue unfolds in data science. After initially selecting a cardiovascular outcome based on precedent in academic literature (piggybacking), “seeing the data, getting to work with data” led them to swap to a renal outcome that was more predictable given available biomarkers. This example illustrates problem (re)formulation strategies serve as the brushstroke of measurement in data science work. Like a painter with a canvas, data scientists apply a strategy, step back to survey their progress along multiple criteria, then consider a new strategy in response to newly-identified flaws.
@@ 5.2. Design Implications: Reconciling Top-Down Measurement Interventions with Bricolage Practice
Our findings uncover a tension between existing interventions designed to help analysts implement the traditional top-down measurement workflow, and data scientists’ bottom-up bricolage practices. However, this tension is not unique to data science. History reveals similar patterns across domains where top-down planning approaches have failed while bottom-up alternatives succeeded. In urban planning, modernist urban renewal projects that imposed regimented public works projects onto complex social systems often destroyed functioning neighborhoods, while community-based planning approaches showed strong success (Fuller and Moore, 2017; Scott, 2020). In software development, rigid and systematic waterfall planning methodologies gave way to agile approaches that embrace iteration and adaptation (Osorio et al., 2011). Key to the success of these bottom-up alternatives was not wholesale rejection of structure, but rather targeted interventions like mixed-use zoning in urban planning and sprint retrospectives in agile development. These interventions recognized that while existing practices were in need of reform, effective solutions must work within resource constraints and pinpoint specific procedural weaknesses in need of constructive redirection. Similarly, when improving measurement practices in data science, interventions tailored to support data scientists’ existing practices are likely to be more effective than direct translation of rigid, top-down measurement frameworks. In the following sub-sections, we examine how existing top-down measurement interventions might be adapted to better support data scientists’ bricolage practices.
@@ 5.3.3. Helping Data Scientists Apply Strategies
Refining. While refining, data scientists make granular adjustments to how a target variable is coded in their data. Participants often engaged in refining by modifying a label definition used to operationalize a construct using a set of indicators used to categorize instances. For instance, P13 engaged in refining by working with domain experts to spot gaps in label definitions, then patch them with a revised set of indicators. Though refining can improve validity, practitioners lack tools to encode domain experts’ knowledge into label definitions. Instead, existing tools developed by the HCI and ML communities help annotators apply a fixed label definition to annotate data (e.g., (Nakayama et al., 2018; Muller et al., 2021; Desmond et al., 2021; Daudert, 2020)) or use crowdsourcing to refine annotation guidelines (e.g., (Chen and Zhang, 2023; Bragg et al., 2018; Chang et al., 2017; K. Chaithanya Manam et al., 2019; Manam and Quinn, 2018)). [^4] Traditional annotation guidelines are de facto label definitions if they serve as a measurement instrument that operationalizes a latent construct of interest. Future HCI and ML research can help data scientists encode domain experts’ knowledge in label definitions by developing tools that help teams collaboratively systematize concepts (Adcock and Collier, 2001) by iteratively identifying and refining indicators used in label definitions. Because our results indicate that domain experts require rich contextual information to identify areas where the systemization of a concept is incomplete, tools supporting this process should capture rich contextual information (e.g., written notes) in data.